Solve for x to the nearest 10th.


63=1.08^x+20



1
Expert's answer
2021-01-26T04:10:04-0500

"63=1.08^x+20"

"43 = 1.08^x"

"log_{1.08}43 = x"

"x = \large\frac{lg43}{lg1.08}" = 48.871

to the nearest tenth is 48.9